### v4

# Router-Tuning: A Simple and Effective Approach for Dynamic Depth

Abstract
The Mixture of Depths (MoD) was introduced to dynamically adjust the computational depth by skipping less important layers, thereby reducing computation and improving efficiency. Despite its promise, existing MoD approaches remain under-explored and face two main challenges: (1) high training costs due to the need to train the entire model along with the routers that determine which layers to skip, and (2) the risk of performance degradation when important layers are bypassed. In response to the first issue, we propose Router-Tuning, which fine-tunes only the routers on a small dataset, drastically reducing the computational overhead associated with full model training. For the second challenge, we investigate Router-Tuning across different architectures and granularities, demonstrating its effectiveness on Attention layers and MoE layers. This method preserves the model’s performance while significantly enhancing computational and memory efficiency. Extensive experiments demonstrate that our approach delivers competitive results while dramatically improving the computation efficiency, e.g., 21% speedup and only a 0.2% performance drop. The code is released at https://github.com/CASE-Lab-UMD/Router-Tuning.
Router-Tuning: A Simple and Effective Approach for Dynamic Depth
Shwai He1 Tao Ge2 Guoheng Sun1 Bowei Tian1 Xiaoyang Wang2 Dong Yu2 1University of Maryland, College Park 2Tencent AI Lab, Bellevue, WA shwaihe@umd.edu
1 Introduction
Large Language Models (LLMs) have shown promising performance across various domains OpenAI et al. (2024); Team (2024); DeepSeek-AI et al. (2024). However, the continuous increase in model size leads to substantial computational costs in real-world applications, making computation reduction a critical research focus for improving LLM efficiency Sun et al. (2024); Lin et al. (2024). A promising approach to this challenge is the Mixture of Depths (MoD) Raposo et al. (2024), which dynamically allocates computational resources for specific inputs. Instead of uniformly applying all layers to every input, MoD selectively activates only a subset of the model’s layers, skipping those deemed less important. This targeted activation significantly reduces computational overhead while maintaining performance.
Despite its potential, current MoD methods are still underexplored and face several critical challenges. On the one hand, the involvement of additional router networks, which decide which layers to skip, often requires extra extensive training: Raposo et al. (2024) train the entire model from scratch while Tan et al. (2024) performs costly continual training. This creates a significant barrier to efficiently integrating MoD with existing LLMs. Furthermore, most prior MoD implementations Raposo et al. (2024); Tan et al. (2024) have been applied to transformer blocks and MLP layers, which are sensitive to skipping. As a result, omitting important components often leads to significant performance degradation He et al. (2024b).
These challenges prompt us to consider two key questions: (1) How can we implement dynamic depth while minimizing training costs and time? (2) How can we enhance model efficiency without compromising performance, achieving faster inference while maintaining competitive task accuracy?
To tackle the first challenge, we introduce Router-Tuning, a novel method that fine-tunes only the router network without updating the backbone model’s parameters. As each router network is a lightweight, single-layer projector that accounts for less than 0.01% of the total parameters, the training overhead is minimal and even significantly lower than that of parameter-efficient finetuning methods Houlsby et al. (2019a); He et al. (2021) like LoRA Hu et al. (2022). Router-tuning requires only a small-scale dataset and fewer training steps, eliminating the need for large-scale pretraining or extensive continual training. This makes it a highly efficient and scalable solution.
To address the second challenge, we conduct a comprehensive investigation of target modules (e.g., Block, MLP, Attention, MoE), and various granularities (e.g., token and sequence). For dense transformer architectures, we propose Attention with Dynamic Depths, which selectively applies dynamic depth to attention layers. By focusing on attention layers known to exhibit high redundancy He et al. (2024b), Router-Tuning not only preserves model accuracy but also alleviates computational and memory bottlenecks. For MoE layers, which significantly impact efficiency due to the computational demands of multiple expert networks, we deploy Router-Tuning on each expert network to enhance the overall efficiency.
Through extensive experiments, we demonstrate the effectiveness of our approach across multiple open-source language models, including Llama Touvron et al. (2023), Mistral Jiang et al. (2023), Qwen Bai et al. (2023), Deepseek-MoE Dai et al. (2024), and OLMoE Muennighoff et al. (2024). Router-Tuning requires less than half an hour on an Nvidia RTX A6000, making it 1000 times faster than DLO Tan et al. (2024). Router-Tuning maintains a high percentage of the original model’s performance while significantly reducing memory usage and accelerating inference, achieving, for example, a 21% inference speedup with only a 0.2% performance degradation. Furthermore, Router-Tuning can be seamlessly integrated with LoRA fine-tuning, further enhancing both efficiency and performance.
In short, our contributions are as follows:
  • We introduce Router-Tuning, a method that fine-tunes only the router on a small dataset, addressing the high computational cost of training the entire model with routers. This approach drastically reduces training overhead while maintaining effectiveness.
  • To explore the trade-off between performance and efficiency, we investigate target modules, deployment granularities, and diverse architectures, demonstrating the effectiveness of Router-Tuning on Attention and MoE layers.
  • Through comprehensive experiments, our approach demonstrates competitive results and significant efficiency improvements. Additionally, Router-Tuning can be seamlessly integrated with LoRA fine-tuning to further enhance both performance and efficiency.
2 Related Work
Layer Redundancy
While increasing the depth of large language models has demonstrated promising performance across a wide range of tasks OpenAI et al. (2024); Team (2024), it also introduces layer redundancy Gromov et al. (2024); He et al. (2024b), posing efficiency challenges. To address this issue, several approaches have been proposed to reduce model depth Men et al. (2024) while maintaining comparable performance. Surprisingly, removing redundant layers has been shown to preserve performance while significantly reducing memory and computational costs Gromov et al. (2024); He et al. (2024a, b). Specifically, Gromov et al.Gromov et al. (2024) suggest dropping continuous Transformer blocks, and He et al. He et al. (2024b) propose fine-grained layer dropping to further improve the effectiveness of layer reduction. However, these static techniques fail to account for the varying complexity of different input sequences, where excessive layer removal can significantly degrade performance on more complex tasks. Instead of statically removing unimportant layers, our approach focuses on dynamically skipping these layers based on the specific inputs.
Dynamic Depth
Dynamic Depth, which allocates different layers based on the specific input, is an effective technique for accelerating inference while preserving performance Han et al. (2021, 2022). Recent works primarily implement dynamic depth through two key methods: Early-Exit Bae et al. (2023); Elhoushi et al. (2024) and Mixture of Depths (MoD) Raposo et al. (2024). Early-exit strategies terminate computation in later layers once sufficient confidence is achieved, effectively reducing redundant computations. In contrast, MoD offers greater flexibility by dynamically skipping less critical layers, enhancing adaptability and representational capacity. Despite their advantages, both Early-Exit and MoD often involve significant training overhead. For instance, LayerSkip Elhoushi et al. (2024) and MoD Raposo et al. (2024) require training models from scratch or extensive continual training, while Tan et al. Tan et al. (2024) extends pre-trained model training over long schedules to achieve optimal performance. To overcome these limitations, we propose Router-Tuning, an efficient approach to dynamic layer skipping that requires minimal additional offline training, providing a more practical and cost-effective solution.
3 Methodology
Figure 1: Overview of Router-Tuning. Router-Tuning involves a trainable router to determine whether a given layer $\bm{F}(\cdot)$ (e.g., Attention and MLP) would be skipped. Inputs with routing scores lower $\bm{R}(\bm{x})$ than the threshold $\tau$ are skipped, and only the router $\bm{R}$ is trainable in the whole model.
In this section, we first review the challenges associated with deploying Mixture of Depths and then introduce Router-Tuning, addressing the implementation of Mixture of Depths from both design and training perspectives.
3.1 Motivation
The Mixture of Depths (MoD) framework Raposo et al. (2024), which dynamically adjusts layer depth based on input complexity to enhance computational efficiency, was originally designed for integration during the pretraining phase, where transformer models are trained from scratch with MoD-enabled layers. More recently, Tan et al. (2024) applied MoD to pretrained Llama models Touvron et al. (2023) through continual training. While these approaches have demonstrated promising results, deploying MoD remains computationally expensive and time-consuming, posing challenges for scalability and real-world deployment. A more efficient alternative is to apply MoD directly to existing pretrained models, followed by lightweight fine-tuning of a subset of parameters Houlsby et al. (2019b); Hu et al. (2022), significantly reducing both computational costs and training time.
On the other hand, MoD has typically been implemented at the transformer block level. However, skipping entire transformer blocks has shown to be suboptimal to maintain the performance. Inspired by He et al. (2024b), we recognize that each transformer block contains layers of varying importance. Aggressively skipping entire blocks risks omitting critical layers, potentially degrading performance. Instead, skipping fine-grained layers offers a more effective strategy for preserving model accuracy. Moreover, unlike blocks that generally share the same architecture, individual layers impose different computational costs. For instance, in dense transformer models, attention layers are particularly expensive, with computational complexity scaling quadratically with sequence length and additional memory needed for KV cache storage. In contrast, in MoE models, MLP layers hold the majority of the parameters, leading to substantial communication and computation overhead.
These insights motivate us to deploy MoD in a cost-effective manner (i.e., Router-Tuning) while optimizing the balance between performance and efficiency.
3.2 Router-Tuning for Dynamic Depth
In this part, we propose Router-Tuning to address the challenges outlined in Section 3.1. As illustrated in Figure 1, Router-Tuning incorporates an additional trainable router that determines whether to skip the layer. Specifically, Router-Tuning can be deployed in two levels: (1) token-level, where layers are dynamically skipped for individual tokens, and (2) sequence-level, where layers are dynamically skipped for the entire sequence. Given an input $\bm{x}\in\mathbb{R}^{L\times d}$, the router first computes an importance score for the input:
$\bm{R}(\bm{x})_{i}=\begin{cases}\text{GATE}(\bm{x}_{i}),&\text{Token-level}\\ \text{GATE}(\frac{1}{L}\sum_{i=1}^{L}\bm{x}_{i}),&\text{Sequence-level}\\ \end{cases},$ (1)
where $\bm{R}$ is a scoring router that assesses the importance score of the input, GATE is the gating function $\text{GATE}(\bm{x})=\text{Sigmoid}(\bm{W}\bm{x})$. Based on the computed importance scores, we further apply a binarized mask $\bm{M}$ to determine whether to skip a token or an entire sequence:
$\bm{M}=\begin{cases}1,&\text{ if }\bm{R}(\bm{x})\geq\tau\\ 0,&\text{ otherwise }\end{cases},$ (2)
where $\tau$ is the threshold. The score is set to zero for skipped inputs and one for retained inputs, ensuring stable outputs Tan et al. (2024).
To enable a differentiable and trainable binary decision process, we utilize the straight-through estimator (STE) Bengio et al. (2013), which allows gradients to propagate through the binary selection mechanism via $\frac{\partial{\bm{M}}}{\partial{\bm{R}}}=1$. The final output of MoD is then computed as follows:
$\bm{y}={\bm{M}}\odot\bm{F}(\bm{x})+\bm{x},$ (3)
where $\bm{F}$ denotes a given layer and $\bm{y}$ is the output. This formulation ensures that the router is fully trainable through the following gradient calculations:
$\frac{\partial\bm{y}}{\partial\bm{W}}=\frac{\partial\bm{y}}{\partial{\bm{M}}}\frac{\partial{\bm{M}}}{\partial{\bm{R}}}\frac{\partial\bm{R}}{\partial{\bm{W}}}.$ (4)
During inference, without the need for gradient calculations, we further enhance computational efficiency by completely bypassing computations for skipped inputs:
$\bm{y}=\begin{cases}{\bm{F}}(\bm{x})+\bm{x},&\text{ if }\bm{R}(\bm{x})\geq\tau\\ \bm{x},&\text{otherwise}\end{cases}.$ (5)
This dynamic routing mechanism ensures that computation is performed only when necessary, thereby enhancing the computational efficiency.
3.3 Extension to Mixture of Experts
Mixture of Experts (MoE) employs sparse activation, dynamically selecting expert networks for each input, which delivers promising performance in various tasks Jiang et al. (2024); Dai et al. (2024); Muennighoff et al. (2024). However, MoE also exhibits significant redundancy, allowing certain experts or layers to be skipped with minimal impact on performance Lu et al. (2024); He et al. (2024a). Building on this, we extend Router-Tuning to MoE layers by implementing dynamic skipping within each expert:
$\hat{\bm{E}}_{i}(\bm{x})=\begin{cases}\bm{E}_{i}(\bm{x}),&\text{ if }\bm{R}(\bm{x})\geq\tau\\ 0,&\text{otherwise}\end{cases},$ (6)
where $\bm{E}_{i}$ denotes the $i$-th expert and $\hat{\bm{E}}_{i}(\bm{x})$ is the corresponding output denotes the corresponding output, bypassing the skipped tokens. Given a collection of $n$ experts, $\{\bm{E}_{1},\bm{E}_{2},\dots,\bm{E}_{n}\}$, the overall output of the MoE layer is as follows:
$\mathcal{K}=\mathrm{TopK}(\mathrm{Softmax}(\bm{G}(\bm{x})),k),$ (7)
$\bm{y}=\sum\nolimits_{i\in\mathcal{K}}{\bm{G}(\bm{x})_{i}\cdot\hat{\bm{E}}_{i}(\bm{x})},$ (8)
where $\mathcal{K}$ denotes the indices of the top-$k$ selected experts, and $\bm{G}(\bm{x})_{i}$ represents the selection score for the $i$-th expert. By dynamically skipping experts within each layer, Router-Tuning significantly reduces both communication and computation costs.
3.4 Training Objectives
Given the computationally intensive nature of training entire LLMs and the constraints of real-world computational resources, our goal is to implement dynamic depth while minimizing both computational costs and time overhead. To achieve this, we focus exclusively on fine-tuning the router, as illustrated in Figure 1, thereby eliminating the need for costly full-model training.
Specifically, we optimize two training objectives: improving task-specific performance and lowering MoD capacity (the proportion of non-skipped inputs). On the one hand, Router-Tuning is designed to maintain the performance of the original model, which we enforce using the loss term $\mathcal{L}_{\text{task}}$ during fine-tuning. On the other hand, the model is encouraged to skip more tokens or sequences (i.e., reduce MoD capacity) to enhance efficiency. To achieve this, we introduce another loss term $\mathcal{L}_{\text{MoD}}$, which drives the model to reduce MoD capacity to a desired target sparsity level $s$, thereby lowering computational costs and accelerating inference. The overall training objective is as follows:
$\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda\cdot\mathcal{L}_{\text{MoD}},$ (9)
$\mathcal{L}_{\text{MoD}}=\text{ReLU}(\|M\|_{0}-s),$ (10)
where $\mathcal{L}$ represents the standard loss function (e.g., cross-entropy), while $\mathcal{L}_{\text{MoD}}$ is an $l_{0}$-norm regularization term that reduces MoD capacity. The coefficient $\lambda$ acts as a scaling factor to balance the trade-off between task performance and efficiency.
4 Experiment Setup
Models
We conduct experiments on Llama Touvron et al. (2023); Grattafiori et al. (2024), Qwen Bai et al. (2023), and Mistral Jiang et al. (2023) due to their competitive performance and widespread adoption. Additionally, we leverage OLMoE Muennighoff et al. (2024) and Deepseek-MoE Dai et al. (2024) as the backbone to deploy Router-Tuning on the Mixture of Experts.
Datasets
For the training dataset, we used Llama-Pro Wu et al. (2024), given it spanning general instruction, math, and code for the SFT process and offering a wealth of instruction data with varying complexity levels. To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 1, which includes multiple tasks: ARC-C Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). The evaluation code is based on EleutherAI’s LM Harness framework Gao et al. (2023).
Table 1: Experimental settings for evaluation tasks. “Norm” refers to the normalization performed with respect to the length of the input.
 Task  Number of few-shot  Metric
BoolQ 0 Accuracy
RTE 0 Accuracy
OBQA 0 Accuracy (Norm)
PIQA 0 Accuracy (Norm)
MMLU 5 Accuracy
WinoGrande 5 Accuracy
GSM8K 5 Exact Match
HellaSwag 10 Accuracy (Norm)
ARC-C 25 Accuracy (Norm)
Hyperparameters
We set $\tau$ as 0.5, which corresponds to the midpoint of the sigmoid function. To ensure that training starts from dense models, we initialize $\bm{W}$ to zero, ensuring that $\bm{R}(\bm{x})\geq\tau$ initially, i.e., training from dense models. To achieve the desired MoD capacity, we perform a grid search over the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} and the scale factor $\lambda$ from {0, 0.1, 0.01, 0.001}, respectively
Table 2: Experimental results of Router-Tuning deployed at different granularities, assessing both task performance and inference efficiency. While Router-Tuning is primarily applied to Attention layers, we also evaluate its performance on Block and MLP layers for comparison. The number of skippable layers is constrained to 16, and the overall MoD capacity is 50%.
Llama-3-8B
Method Granularity Speedup ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline     – 1.00×1.$0$ 58.1 81.3 82.1 65.3 45.0 80.5 67.2 77.7 69.7
Router-Tuning $\mathrm{Block}_{token}$ $\bf 1.24\times$ 44.2 77.9 63.1 64.4 34.0 70.4 65.4 71.6 61.4
$\mathrm{Block}_{seq}$ $\bf 1.26\times$ 44.5 78.0 62.6 64.6 34.2 70.3 65.3 71.2 61.3
$\mathrm{MLP}_{token}$ 1.05×1.05$\times$1.05 × 45.3 77.8 65.1 62.8 33.7 71.9 66.8 72.4 62.0
$\mathrm{MLP}_{seq}$ 1.06×1.06$\times$1.06 × 45.1 77.7 65.4 62.4 33.4 71.6 66.4 72.1 61.8
$\mathrm{Attn}_{token}$ $\bf 1.18\times$ 56.4 79.8 81.0 65.3 45.2 79.9 64.6 77.3 68.7
$\mathrm{Attn}_{seq}$ $\bf 1.21\times$ 56.6 80.5 80.7 65.1 44.6 80.5 69.7 77.7 69.4
Llama-3-8B-Instruct
Method Granularity Speedup ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline     – 1.00×1.$0$ 62.1 83.2 78.8 65.7 42.8 78.7 67.5 75.9 69.3
Router-Tuning $\mathrm{Block}_{token}$ $\bf 1.24\times$ 44.6 80.9 54.1 60.2 31.2 64.8 67.7 65.1 58.6
$\mathrm{Block}_{seq}$ $\bf 1.26\times$ 44.7 81.2 54.5 60.6 32.4 64.6 67.1 64.8 58.7
$\mathrm{MLP}_{token}$ 1.05×1.05$\times$1.05 × 41.4 74.9 59.3 64.8 31.6 67.8 66.4 68.4 59.3
$\mathrm{MLP}_{seq}$ 1.06×1.06$\times$1.06 × 41.8 75.1 59.3 64.5 31.2 68.2 66.7 68.8 59.5
$\mathrm{Attn}_{token}$ $\bf 1.18\times$ 60.2 82.9 76.8 65.8 42.6 78.6 67.7 76.6 68.9
$\mathrm{Attn}_{seq}$ $\bf 1.21\times$ 60.4 83.3 76.9 65.7 43.0 78.2 68.2 76.9 69.1
5 Main Results
In this section, we evaluate the effectiveness of Router-Tuning on transformer architectures with the deployment details in Appendix 4.
5.1 Performance of Router-Tuning
Router-Tuning achieves superior performance in Attention layers We first compare deploying Router-Tuning to different modules, e.g., Block, MLP, and Attention, as shown in Table 2. Based on the observation that deeper layers are more redundant than shallow layers Gromov et al. (2024); He et al. (2024b), we focus on deploying Router-Tuning to the deepest layers except the last one, leaving other layers unchanged.
While previous studies have primarily explored layer dropping or skipping to Block and MLP layers Bae et al. (2023); Gromov et al. (2024), skipping these modules significantly degrades performance when applied at either token or sequence level. In contrast, applying dynamic depth to Attention layers maintains the performance of original models, e.g., 69.4 v.s. 69.7 in Llama-3-8B. These findings reinforce our motivation to target Attention layers, and we utilize Router-Tuning on Attention layers by default unless stated otherwise.
Figure 2: Comparison with Attention Drop under identical skipping ratios, highlighting the performance benefits of Router-Tuning.
Figure 3: Layer-wise Skipping Ratios for MoD and Attention Drop under overall skipping ratios of $50\%$.
Router-Tuning significantly outperforms static layer dropping
While Attention Drop He et al. (2024b), which statically drops attention layers, has demonstrated the promising performance of Layer Drop, its static nature lacks flexibility and limits representational power. Here, we further investigate the improvements offered by the dynamic mechanism. As shown in Figure 2, we compare Router-Tuning with Attention Drop He et al. (2024b) under the same computation budget — for instance, dropping 4 layers versus deploying MoD to 8 layers with 50% capacity. Under the same skipping ratios, Router-Tuning significantly outperforms Attention Drop He et al. (2024b) on the GSM8K benchmark Cobbe et al. (2021), e.g. 6.5% when the skipping ratio is 25.0%. In Figure 3, we further visualize the layer-wise skipping ratios of MoD versus Attention Drop. Unlike static approaches that permanently remove certain layers, Router-Tuning maintains the utilization of all layers by adaptively distributing skipping ratios across them. This flexible allocation strategy contributes to improved performance.
Table 3: Ablation study of Router-Tuning across different models architectures, including Llama Touvron et al. (2023), Qwen Bai et al. (2023), Mistral Jiang et al. (2023).
Models Speedup  OBQA  PIQA  RTE  WinoGrande  BoolQ  ARC-C  HellaSwag  MMLU   Avg.
Llama-2-13B 1.00×1.$0$ 45.2 80.5 65.0 76.2 80.7 59.4 82.2 54.6 68.0
w/Router-Tuning $\bf 1.22\times$ 45.4 80.6 64.6 76.2 80.5 59.3 82.2 54.7 67.9
Qwen-2.5-14B 1.00×1.$0$ 45.6 82.2 79.1 80.4 85.3 67.2 84.3 79.7 75.5
w/Router-Tuning $\bf 1.18\times$ 45.4 82.4 77.9 78.5 85.0 66.2 83.8 78.0 74.7
Qwen-2.5-7B 1.00×1.$0$ 47.2 79.6 81.2 76.6 84.6 63.7 80.2 74.1 73.4
w/Router-Tuning $\bf 1.19\times$ 47.0 80.1 76.9 76.1 83.2 62.3 79.8 73.3 72.3
Mistral-7B 1.00×1.$0$ 44.4 82.2 68.2 79.0 82.2 60.6 83.2 62.4 70.3
w/Router-Tuning $\bf 1.24\times$ 44.0 81.8 67.6 78.2 81.7 59.9 82.6 61.8 69.7
5.2 Efficiency Improvements
In this part, we measure the efficiency in both training and inference, focusing on computational and memory usage.
Training Efficiency
The training efficiency of our method lies in two perspectives: trainable parameters and training steps. Since the router projects the input from dimension $d$ to 1, the number of trainable parameters is $d\times 1$ per layer, and the total number of trainable parameters is fewer than 0.01% of the whole model. Additionally, Router-Tuning only requires a few steps, which is verified in Section 6. Consequently, Router-Tuning on a single Nvidia A6000 GPU only takes less than 15 minutes, which is over 1000 times faster than DLO Tan et al. (2024) that finetunes the entire model and requires 36 hours on Nvidia RTX A100 GPUs.
Inference Speedup
We also evaluate the run-time speed improvements achieved with Router-Tuning. The inference speed is measured throughout the entire generation process, including prefilling and generation. To ensure that our results accurately reflect the performance gains, we adhere to two key principles: (1) all operations are performed on a single Nvidia RTX A6000 Ada GPU, eliminating any communication overhead from multi-GPU setups; and (2) we increase batch sizes to fully utilize the GPU for each model. As shown in Table 2, different granularities contribute to different levels of speedup. This variation is primarily due to attention layers, where fine-grained token-level MoD introduces differing token lengths within a batch, necessitating padding operations to standardize sequence lengths. Instead, the speedups at the sequence level surpass those at the token level, with Router-Tuning achieving a 21% improvement in inference speed.
KV Cache
The KV cache stores intermediate representations of attention layers, accelerating inference by eliminating redundant computations but incurring substantial memory overhead. Our approach, which selectively skips attention layers, significantly reduces KV cache size—for instance, achieving an 8GB reduction when processing an input sequence of length 2048 with a batch size of 64 on Llama-3-8B. In contrast, DLO Tan et al. (2024) operates exclusively on MLP layers and retains the full KV cache, providing no memory savings.
Figure 4: Ablation study on the number of MoD layers.
6 Ablation Studies
We conduct ablation studies to evaluate the generality and robustness of Router-Tuning across various settings, including different model architectures, the number of MoD layers, training datasets.
Table 4: Performance of Router-Tuning on Mixture of Experts, where we take Expert Drop He et al. (2024a) that statically skips the experts as the baseline of static dropping for comparison.
Models SpeedUp OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag  MMLU   Avg.
DeepSeek-MoE 1.00×1.$0$ 43.6 80.5 62.8 73.4 72.4 52.7 79.9 44.5 63.7
w/Drop $\bf 1.11\times$ 42.2 80.2 59.9 70.0 74.0 48.1 75.6 38.9 61.1
w/Router-Tuning $\bf 1.10\times$ 43.2 80.4 61.2 71.4 72.1 50.8 77.3 42.8 62.4
OLMoE 1.00×1.$0$ 45.6 80.1 53.7 71.2 74.7 54.5 79.4 52.5 64.0
w/Drop $\bf 1.13\times$ 34.0 66.6 51.6 59.3 67.6 39.0 40.5 42.7 50.2
w/Router-Tuning $\bf 1.12\times$ 40.4 76.2 53.2 70.2 71.3 52.0 77.9 50.1 61.4
Compatibility across different models
Since Router-Tuning can be seamlessly integrated into pretrained language models, we extend our evaluation to diverse architectures, including Llama-2 Touvron et al. (2023), Mistral Jiang et al. (2023), and Qwen2.5 Bai et al. (2023), covering a wide range of model sizes. As shown in Table 3, we deploy Router-Tuning in half of the attention layers while maintaining the total MoD capacity at 50%. Across all models, Router-Tuning preserves performance compared to their original counterparts, demonstrating its effectiveness and adaptability across different architectures.
Figure 5: Ablation study on training samples.
Impact of number of MoD layers
In the main experiments, we deployed half of the layers with MoD. In Figure 4, we further explore the effect of the number of MoD layers. Our results indicate that applying MoD to up to half of the attention layers still maintain comparable performance. However, when further increasing the number of MoD layers, performance starts to degrade. We attribute this decline to the transformation of important shallow layers, which negatively impacts overall performance Men et al. (2024); He et al. (2024a, b). Therefore, preserving the density of shallow layers while applying MoD to deeper layers ensures its effectiveness.
Influence of Training Dataset
In Table 5, we next examine the impact of using different training datasets for Router-Tuning. We consider a variety of datasets, including Alpaca Taori et al. (2023), Evol-Instruct Xu et al. (2023), ShareGPT Zheng et al. (2023), and Llama-Pro Wu et al. (2024). Since Router-Tuning only fine-tunes the routers while keeping the backbone of the language models intact, changes in the training dataset do not significantly impact performance. However, Llama-Pro, which incorporates diverse training data from various domains, provides slightly better performance due to its broader data coverage.
Table 5: Effectiveness across different training datasets, where Router Tuning demonstrates robustness to the varying datasets.
Dataset HellaSwag  MMLU  OBQA  WinoGrande Avg.
Baseline 82.1 65.3 45.0 77.7 67.5
Alpaca 79.8 62.2 43.8 77.4 65.8
Evol-Instruct 80.4 64.0 44.4 77.6 66.6
ShareGPT 80.6 63.3 45.4 76.7 66.5
Llama-Pro 80.7 65.1 44.6 77.7 67.0
On the other hand, due to the small number of trainable parameters, Router-Tuning does not require a large amount of training samples. As shown in Figure 5, MoD layers are initially dense-activated and then sparsified. Although the initial sparsification steps lead to a drop in performance, subsequent Router-Tuning facilitates performance recovery. Notably, just 5K training samples are sufficient to effectively train the routers.
7 Router-Tuning Extensions
In this section, we explore the extensions of Router-Tuning to broader architectural designs and training paradigms. Specifically, we demonstrate its applicability to mainstream Mixture-of-Experts (MoE) architectures, and examine its integration with LoRA-based fine-tuning to jointly improve efficiency and downstream performance.
Table 6: Effectiveness of Router-Tuning integrated with LoRA fine-tuning, compared to using Router-Tuning and LoRA independently. The integrated approach achieves comparable performance while preserving efficiency, highlighting the compatibility and complementary strengths of the two methods.
Model Method   OBQA PIQA   RTE WinoGrande   BoolQ ARC-C   HellaSwag MMLU   Avg.
Llama-3-8B Baseline 45.0 80.5 67.2 77.7 81.3 58.1 82.1 65.3 69.6
Router-Tuning 44.6 80.5 69.7 77.7 80.7 56.6 80.7 65.1 69.5
LoRA 46.6 82.0 68.0 77.9 83.9 61.8 81.6 65.9 71.0
   LoRA + Router-Tuning 47.2 82.2 67.4 77.8 83.9 61.5 81.7 65.8 70.9
Mistral-7B Baseline 44.4 82.2 68.2 79.1 82.2 60.6 83.2 62.4 70.3
Router-Tuning 44.2 81.9 68.5 78.6 81.7 60.4 82.5 61.8 70.0
LoRA 45.2 83.0 68.9 79.4 84.7 60.7 83.7 62.8 71.1
LoRA + Router-Tuning 45.7 83.1 68.7 79.3 84.3 60.9 83.4 62.9 71.2
Router-Tuning on MoE
The Expert’s redundancy in MoE has been widely demonstrated in recent works Lu et al. (2024); He et al. (2024a), e.g., the model still maintains comparable performance after removing certain layers. Therefore, we further extend Router-Tuning to MoE, where we take OLMoE Muennighoff et al. (2024) and DeepSeek-MoE Dai et al. (2024) as the backbones and equip each Expert network with Router-Tuning. Since these models deploy MoE at the token level, where each expert processes tokens from different sequences, we directly apply token-level Router-Tuning to these models. Here, we compare Router-Tuning with Expert Drop He et al. (2024a) that drops less important experts. For deployment, we configure an overall skipping ratio of 25%, achieved either by statically dropping 25% of layers or utilizing MoD on half of the MoE layers with 50% MoD capacity.
Figure 6: Expert-wise MoD capacity, visualizing the capacity utilization and selected frequency of individual experts within a specific layer.
In Table 4, while Expert Drop affects the performance significantly, Router-Tuning preserves more than 95% of the original performance, which demonstrates Router-Tuning’s effectiveness in the MoE architecture. In Figure 6, we further visualize the capacity utilization of individual experts within a specific layer, considering two perspectives for each expert: the frequency of expert selection (“Frequency”) and the MoD capacity allocated to each expert (“Capacity”).
Instead of removing a subset of experts (typically those being selected less frequently), Router-Tuning adopts a specialized and efficient utilization pattern, where high-frequency experts undergo greater sparsification compared to low-frequency experts. This approach retains the potential of all experts and fosters a more balanced utilization across the entire set of experts Fedus et al. (2022). Therefore, this approach maintains the potential of all experts and promotes a more balanced utilization across all experts.
Integration with LoRA Router-Tuning finetunes only a small subset of parameters, sharing similarities with parameter-efficient finetuning (PEFT) in terms of reduced parameter updates but differing in objectives: Router-Tuning focuses on achieving dynamic depth to enhance efficiency, whereas PEFT aims to adjust a small number of parameters to improve downstream task performance. Given that Router-Tuning is orthogonal to PEFT, we propose integrating Router-Tuning with LoRA finetuning Hu et al. (2021) to simultaneously enhance both task performance and efficiency. As the results shown in Table 6, jointly training routers and LoRA also preserves the performance of LoRA finetuning. Therefore, this joint training strategy serves as an advanced setting that further enhances the overall model capability.
8 Conclusion
In this work, we investigate the dynamic depth mechanism from both design and training perspectives. We propose Router-Tuning, which effectively enhances computational efficiency while preserving model performance. Furthermore, Router-Tuning implements dynamic depth by fine-tuning only a minimal number of parameters in just a few steps. Additionally, we explore Router-Tuning across a variety of modules and granularities to evaluate its effectiveness across a wide range of models and tasks. These advancements provide valuable insights and practical solutions for improving the efficiency of large language models.
Limitations
Despite the progress we have made, our work still has limitations. First, while we have advanced MoD with Router-Tuning, other sophisticated training methods may exist that could further improve performance, warranting future exploration. Second, due to computational resource constraints, our experiments have been limited to a small set of models and tasks. Expanding this approach to other models and a broader range of tasks would be highly valuable for understanding its full potential.